OpenStudy (anonymous):

The half-life of a certain radioactive material is 32 days. An initial amount of the material has a mass of 361 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 5 days. Round your answer to the nearest thousandth.

OpenStudy (phi):

do you know a formula for half-life? see https://en.wikipedia.org/wiki/Half-life#Formulas_for_half-life_in_exponential_decay

OpenStudy (anonymous):

Idk how to actually do it

OpenStudy (phi):

I would use this formula \[ Amt= M \left(\frac{1}{2} \right)^{\frac{t}{t_0}} \] in your problem, M is 361 kg t0 is 32 days

OpenStudy (johnweldon1993):

Indeed^ and for the quetion, to find out how much would remain after 5 days, you would make 't' = 5 so \[\huge \text{Amount} = 361(\frac{1}{2})^{\frac{5}{32}}\]

OpenStudy (anonymous):

Sooo?? im lost

OpenStudy (johnweldon1993):

The equation that phi has given is exactly what the first half of the question asks, "write an exponential function for this decay" So that would be just plugging in M and t_o into the equation for half life \[\large \text{Amount} = 361(\frac{1}{2})^\frac{t}{32}\] This would be the general equation, Now to find the amount after 5 days, you plug in 5 for 't' \[\large \text{Amount} = 361(\frac{1}{2})^\frac{5}{32}\] Now you just plug that into a calculator and solve

OpenStudy (anonymous):

28.203125 ?

OpenStudy (anonymous):

Or is that completely wrong ?

OpenStudy (johnweldon1993):

Completely wrong, but that's alright, probably just missed a parenthesis 361(1/2)^(5/32) is what you would plug into a calculator

OpenStudy (anonymous):

@$w3G_Godd

OpenStudy (anonymous):

is the question multiple choice

OpenStudy (anonymous):

because i remember having this question but it was multiple choice

OpenStudy (anonymous):

Yeah but they are going to take me forever to type them

OpenStudy (anonymous):

i still have where i did it what does your letter B say

OpenStudy (anonymous):

the answer is here http://www.wolframalpha.com/input/?i=361*%28.5%29^%285%2F32%29

OpenStudy (anonymous):

\[\huge 323.945\]

OpenStudy (anonymous):

Oh okay thanks :)

OpenStudy (anonymous):

Its number 3 on here D:

OpenStudy (anonymous):

but ya lol hes right